# Particle motion and tidal force in a non-vacuum-charged naked singularity

### v1

@@ Abstract
In this article, we investigate the gravitational field of a charged, non-vacuum, non-rotating, spherically symmetric body of mass $M$ assuming a static solution to the Einstein-Maxwell field equations. We demonstrate that the characteristics of perihelion precession of orbits in Schwarzschild and Null singularity spacetimes are similar, whereas, in the case of charged naked singularity (CNS) spacetime, the characteristics are the opposite of those in Schwarzschild and Null singularity spacetime. Here we discuss some novel features of the timelike and lightlike geodesics in this spacetime. We also discuss the comparative study of tidal force in the null singularity spacetime and charged naked singularity spacetime.
@@ I Introduction
The Event Horizon Telescope (EHT) collaboration, by releasing the first image of the astrophysical black hole at the center of M87 galactic center event2019first ; akiyama20196first and the subsequent image of the black hole at the center of our own galaxy - Sagittarius A* (Sgr A*) akiyama2022first has opened up new horizons in the field of astrophysics and black hole physics. Theoretically, there are many possible solutions of Einstein’s field equations other than black hole (BH). A naked singularity (NS) is one such example PhysRevLett.20.878 ; perlick . Null geodesics can escape from such a singularity (i.e., NS) and would be observable to a distant observer under certain conditions of gravitational collapse Joshi:2011zm ; Joshi:2013dva . The singularity could then be locally or globally visible depending upon the time of formation of trapped surfaces and the apparent horizon. Recent observational studies of modified gravity, suggest alternative models of the compact objectKhodadi:2022pqh ; Khodadi:2021gbc ; Khodadi:2020gns ; KumarWalia:2022ddq . In the case of NS, most of the studies are focused around the accretion discs Tahelyani:2022uxw ; Patra:2023epx ; Kovacs:2010xm , dark matter Joshi:2022azj , shadows cast by them Kumar:2020ltt ; Shaikh:2022ivr ; Ghosh:2021txu , and particle trajectories around them Bambhaniya:2019pbr .
@@
Further, it is generally believed that a test body approaching a static, spherically symmetric compact object experiences compression in the angular and spaghettification in the radial directions, as in the case of Schwazchild BH and FNS Madan:2022spd . In the case of electrically charged compact objects (e.g., RN BH) it has been shown by Crispino et al., Crispino:2016pnv that radial and angular components of the tidal effects experienced by a freely falling test body changes sign at the null hypersurface. It is interesting to note that the charge-to-mass ratio of the charged compact object and the position of the test body determine whether it would experience stretching or compression in a radial or angular direction in the vicinity of a charged compact object. The above facts motivated us to make a comparative study on the effects of tidal forces experienced by a test body in the vicinity of FNS and RN spacetimes in our present work.
@@ II Charged Naked Singularity Spacetime (CNS)
where $d\Omega^{2}=d\theta^{2}+sin^{2}\theta d\phi^{2}$ and $M$ is the ADM mass of the spacetime. It is shown in  joshi2020shadow that there is a strong null singularity at the center of the spacetime in Eq. (1). There is no event horizon around the singularity. Moreover, this future null singularity (FNS) spacetime satisfies both weak and strong energy conditions  joshi2020shadow . For our present investigation, we introduce an electric charge in FNS spacetime, and the corresponding action for the electromagnetic field can be expressed as:
@@
This corresponds to a static spherically symmetric body where the charged matter is distributed over the spacetime. The static charge that surrounds such an object contributes to the energy distribution of the spacetime, modifying the spacetime structure. For example, when a (electric) charge is introduced in a Schwarzschild BH, resulting in the RN BH, the nature of singularity changes, i.e., from spacelike to timelike. However, there is a large possibility that the charge near the singularity plays a significant role in the instability of spacetime singularity. In the CNS case, as shown in Fig. (1), spacetime singularity is timelike. Here we do not claim that this singularity is stable instead we show that charge in FNS spacetime changes a causal structure similar to the RN case. The corresponding Penrose diagram for the given metric (9) is in the Fig. (1).
@@
Presented metric (9) is asymptotically flat (Minkowskian) and the inclusion of (electric) charge in this spacetime is attributed to the form of the chosen action, similar to that of RN black hole spacetime. The corresponding Kretschmann scalar ($K=R_{\alpha\beta\gamma\delta}\,R^{\alpha\beta\gamma\delta}$) and Ricci scalar for this new Charged Naked Singularity (CNS) are:
@@ III Timelike and Lightlike geodesics in Charged naked singularity spacetimes
As discussed in II, CNS spacetime is asymptotically flat and harbors a central strong curvature singularity which is not covered by an event horizon. It has temporal and rotational (about the azimuthal angle) symmetries leading to two associated Killing vectors: $\xi^{\mu}_{t}=(1,0,0,0)$ and $\xi^{\mu}_{\phi}=(0,0,0,1)$, such that $\xi^{\mu}_{t}\,\partial_{\mu}=\partial_{t}$ and $\quad\xi^{\mu}_{\phi}\,\partial_{\mu}=\partial_{\phi}$. The corresponding conserved quantities:
@@ IV.3 Comparison of Tidal force between CNS, RN, FNS, and Schwarzchild spacetimes
When we compare the nature of the tidal forces in CNS with Schwarzchild, RN, and FNS spacetimes, we find some distinct features. We observe that for the radial component of the tidal forces, there is an infinite compression for CNS similar to RN spacetime, and in contrast to Schwarzchild and FNS spacetimes. It is to be noted that in CNS, there would be small spaghettification followed by infinite compression for smaller values of $Q$ (i.e., for $Q$ slightly greater than $M$). As we increase the value of $Q$, spaghettification decreases and for large values of $Q$ there would be no spaghettification (in contrast to FNS and Schwarzchild spacetimes). In RN spacetime, there is a rapid change in spaghettification as we increase the value of $Q$ and for large values of $Q$, there would be no spaghettification, similar to CNS. In FNS, there is a finite spaghettification followed by a finite compression while there is infinite spaghettification and no compression in Schwarzchild spacetime as the test body approaches the spacetime singularity. For all the three cases, $Q>M\quad Q=M$ and $Q<M$, the compression sets-in much earlier w.r.t. RN and FNS spacetimes
@@ V Results and Discussion
      The Figures [4 - 5] show the behaviour of the radial and angular tidal forces for different values of charge $Q$ when mass $M$ is kept fixed at $1$. It is to be noted that for radial tidal forces, there is infinite compression for CNS while for the angular part there is infinite spaghettification, respectively.
@@
      Figures (6 - 6) show a comparison of radial tidal force while figures (6 - 6) show the comparison for the angular tidal force for CNS, Schwarzchild, RN and FNS spacetimes. It is observed that both spaghettification and compression set-in earlier in CNS as compared to the RN spacetime.

### v2

@@ Abstract
We investigate the gravitational field of a charged, non-vacuum, non-rotating, spherically symmetric body of mass $M$ assuming a static solution to the Einstein-Maxwell field equations. We show the characteristics of perihelion precession of orbits in the case of charged naked singularity (CNS) spacetime. Here we discuss some novel features of light-like geodesics in this spacetime. We also discuss the comparative study of tidal force in the null singularity spacetime and charged naked singularity spacetime.
@@ I Introduction
The Event Horizon Telescope (EHT) collaboration, by releasing the first image of the astrophysical black hole at the center of M87 galactic center event2019first ; akiyama20196first and the subsequent image of the black hole at the center of our own galaxy - Sagittarius A* (Sgr A*) akiyama2022first has opened up new horizons in the field of astrophysics and black hole physics. Theoretically, there are many possible solutions of Einstein’s field equations other than a black hole (BH). A naked singularity (NS) is one such example PhysRevLett.20.878 ; perlick . Null geodesics can escape from such a singularity (i.e., NS) and would be observable to a distant observer under certain conditions of gravitational collapse Joshi:2011zm ; Joshi:2013dva . The singularity could then be locally or globally visible depending upon the time of formation of trapped surfaces and the apparent horizon. Recent observational studies of modified gravity, suggest alternative models of the compact objectKhodadi:2022pqh ; Khodadi:2021gbc ; Khodadi:2020gns ; KumarWalia:2022ddq . In the case of NS, most of the studies are focused around the accretion discs Tahelyani:2022uxw ; Patra:2023epx ; Kovacs:2010xm , dark matter Joshi:2022azj , shadows cast by them Kumar:2020ltt ; Shaikh:2022ivr ; Ghosh:2021txu , and particle trajectories around them Bambhaniya:2019pbr ; Battista:2022krl .
@@
Further, it is generally believed that a test body approaching a static, spherically symmetric compact object experiences compression in the angular and spaghettification in the radial directions, as in the case of Schwazchild BH and FNS Madan:2022spd . In the case of electrically charged compact objects (e.g., RN BH) it has been shown by Crispino et al., Crispino:2016pnv ; Arora:2023ltv that radial and angular components of the tidal effects experienced by a freely falling test body changes sign at the null hypersurface. It is interesting to note that the charge-to-mass ratio of the charged compact object and the position of the test body determine whether it would experience stretching or compression in a radial or angular direction in the vicinity of a charged compact object. The above facts motivated us to make a comparative study on the effects of tidal forces experienced by a test body in the vicinity of FNS and RN spacetimes in our present work.
@@ II Charged Naked Singularity Spacetime (CNS)
where $d\Omega^{2}=d\theta^{2}+sin^{2}\theta d\phi^{2}$ and $M$ is the ADM mass of the spacetime. It is shown in  joshi2020shadow that there is a strong null singularity at the center of the spacetime in Eq. (1). The Penrose diagram of spacetime can be used to determine the type of spacetime singularity. The Penrose diagram allows us to explain the entire spacetime manifold in a limited-size causal diagram by transforming the temporal and radial coordinates. The compactification of coordinates is as follows:
$T=\tan^{-1}(t+r^{*})+\tan^{-1}(t-r^{*})$
$R=\tan^{-1}(t+r^{*})-\tan^{-1}(t-r^{*})$ (2)
where for Eq. (1),
$r^{*}=r-\frac{M^{2}}{r}+2M\log r.$ (3)
In order to determine the nature of the singularity at $r\to 0$, we need to check the value of $T$ and $R$ as $r\to 0$. From the above Eq. (3), it can be verified that as $r\to 0$ for any finite value of $t$, $r^{*}\to-\infty$, clearly suggesting that the singularity thus formed at $r=0$ corresponding to the line element defined in Eq. (1) is null in nature. (Similarly one would get a timelike singularity if for any finite $t$, $r^{*}\to 0$ when $r\to 0$) (See e.g., dDey ). There is no event horizon around the singularity. Moreover, this future null singularity (FNS) spacetime satisfies both weak and strong energy conditions  joshi2020shadow . For our present investigation, we introduce an electric charge in FNS spacetime, and the corresponding action for the electromagnetic field can be expressed as:
@@
This corresponds to a static spherically symmetric body where the charged matter is distributed over the spacetime. The static charge that surrounds such an object contributes to the energy distribution of the spacetime, modifying the spacetime structure. For example, when an electric charge is introduced in a Schwarzschild BH, resulting in the RN BH, the nature of singularity changes, i.e., from spacelike to timelike. However, there is a large possibility that the charge near the singularity plays a significant role in the instability of spacetime singularity. In the CNS case, as shown in Fig. (1), spacetime singularity is timelike. Here we do not claim that this singularity is stable instead we show that charge in FNS spacetime changes a causal structure similar to the RN case. The corresponding Penrose diagram for the given metric (11) is in the Fig. (1).
@@
Presented metric (11) is asymptotically flat and the inclusion of electric charge in this spacetime is attributed to the form of the chosen action, similar to that of RN black hole spacetime. The corresponding Kretschmann scalar ($K=R_{\alpha\beta\gamma\delta}\,R^{\alpha\beta\gamma\delta}$) and Ricci scalar for this new Charged Naked Singularity (CNS) are:
@@ III Timelike and Lightlike geodesics in Charged naked singularity spacetimes
As discussed in Sec. II, CNS spacetime is asymptotically flat and harbors a central strong curvature singularity which is not covered by an event horizon. It has temporal and rotational (about the azimuthal angle) symmetries leading to two associated Killing vectors: $\xi^{\mu}_{t}=(1,0,0,0)$ and $\xi^{\mu}_{\phi}=(0,0,0,1)$, such that $\xi^{\mu}_{t}\,\partial_{\mu}=\partial_{t}$ and $\quad\xi^{\mu}_{\phi}\,\partial_{\mu}=\partial_{\phi}$. The corresponding conserved quantities:
@@ IV.3 Comparison of Tidal force between CNS, RN, FNS, and Schwarzchild spacetimes
When we compare the nature of the tidal forces in CNS with Schwarzchild, RN, and FNS spacetimes, we find some distinct features. We observe that for the radial component of the tidal forces, there is an infinite compression for CNS similar to RN spacetime, and in contrast to Schwarzchild and FNS spacetimes. It is to be noted that in CNS, there would be small spaghettification followed by infinite compression for smaller values of $Q$ (i.e., for $Q$ slightly greater than $M$). As we increase the value of $Q$, spaghettification decreases and for large values of $Q$ there would be no spaghettification (in contrast to FNS and Schwarzchild spacetimes). In CNS spacetime, there is a rapid change in spaghettification as we increase the value of $Q$ and for large values of $Q$, there would be no spaghettification, similar to RN. In FNS, there is a finite spaghettification followed by a finite compression while there is infinite spaghettification and no compression in Schwarzchild spacetime as the test body approaches the spacetime singularity. For all the three cases, $Q>M\quad Q=M$ and $Q<M$, the compression sets-in much earlier w.r.t. RN and FNS spacetimes
@@ V Results and Discussion
      The Figures [4 - 6] show the behaviour of the radial and angular tidal forces for different values of charge $Q$ when the mass $M$ is kept fixed at $1$. It is to be noted that for radial tidal forces, there is infinite compression for CNS while for the angular part there is infinite spaghettification, respectively.
@@
      Figures (5 - 5) show a comparison of radial tidal force while figures (5 - 5) show the comparison for the angular tidal force for CNS, Schwarzchild, RN and FNS spacetimes. It is observed that both spaghettification and compression set-in earlier in CNS as compared to the RN spacetime.